Plum Corporation will begin operations on January 1. Earnings for the next five years are projected to be relatively stable at about $80,000 per year. The shareholders of Plum are in the 33% tax bracket and dividends are taxable at 15%. Click here to access the tax rate schedule to use for this problem. If an amount is zero, enter O. When required, round your answers to the nearest dollar. a. Assume that Plum will reinvest its after-tax earnings in the growth of the company. If Plum operates as a C corporation, the corporations income tax will be$ o X, and the shareholders liability will be If Plum operates as an S corporation, the corporations income tax will be $ and the shareholders liability will be Therefore, viewed from an entity-owner perspective, operating as will result in overall tax savings. Feedback b. Assume that Plum will distribute its after-tax earnings each year to its shareholders. If Plum operates as a C corporation, the corporations income tax will be $ , and the shareholders liability will be If Plum operates as an S corporation, the corporations income tax will be $ and the shareholders liability will be
Answer #1

Ans Case a : If Corporation Reinvest it's after tax earnings

1. Operating as C Corporation

a. Earnings for the Year $ 80,000

b. Tax Payable =13,750+34%(Between $75,000 and $1,00,000)

c. Corporation Tax = 13,750 +34% (80,000-75,000) = 13,750+34%(5,000)=15,450

d. Earnings After Tax(a-c)= 80,000-15,450= $ 64,550

e. Dividends to shareholders = Nil

f. Shareholder's Liability = Nil

Note : Since the C Corporation is retaining it's Entire Earnings. Hence no dividend is distributed to shareholders. In case on no dividend distribution the dividend distribution tax will be nil and the shareholder's liability will be nil .

Corporation Income Tax(Point c )=15,450 , Shareholder's Liability = 0

2. Operating as S Corporation

Since the tax rates applicable to C corporation will be same as that in case of S corporation and the

Entity retains it's Entire earnings hence in the given case . Answer will be same as that for C corporation above

  Corporation Income Tax =15,450 , Shareholder's Liability = 0

Case b. : If Corporation distributes it's after tax earnings  

1. Operating as C corporation   

a. Earnings for the Year $ 80,000  

b. Tax Payable =13,750+34%(Between $75,000 and $1,00,000)

c. Corporation Tax = 13,750 +34% (80,000-75,000) = 13,750+34%(5,000)=15,450

d. Earnings After Tax(a-c)= 80,000-15,450= $ 64,550

e. Dividends to be distributed to Shareholders = Earnings after Tax =$ 64,550

f. Tax rate applicable on Dividends = 15%

g. Shareholder's Liability = Tax on Dividends = $64550*15% = $9683(approx)

Note : In given case Entire after tax earnings of C corporation are distributed to shareholders hence in the given case Tax on Dividends are calculated on entire after tax earnings .

Corporation Tax = $15,450 and Shareholder's Liability = $9683(approx)

2. Operating as S corporation

Same answer as applicable to C corporation Corporation Tax = $15,450 and Shareholder's Liability = $9683(approx)

(In the Given Case Since tax rates and after tax earnings distributed are same hence the answer arrived will be same as that for C corporation .)